OpenStudy (anonymous):

Solve for "r": 4/3pi(r^3)=V

OpenStudy (ghazi):

\[\frac{ 4 }{ 3 } \pi r^3=V\] divide both sides by pi and 4/3 after that you'll have \[r^3=\frac{ 3 }{ 4 }V/ \pi\]

OpenStudy (ghazi):

finally \[r=\sqrt[3]{\frac{ 3 }{ 4 } V/ \pi}\]

OpenStudy (anonymous):

That's kind of what I thought but the problem itself looked really scary. Thank you for your help!

OpenStudy (ghazi):

:) YW